10^(10x-7)=3^(2x-4)
Answer by lwsshak3(11628)

solve for x
10^(10x-7)=3^(2x-4)
take log of both sides
(10x-7)log10=(2x-4)log3
(10x-7)/(2x-4)=log3=.4771
(10x-7)=(2x-4)*.4771
10x-7=.9542x-1.9085
9.0458x=5.0915
x=5.0915/9.0458=.5629